OpenStudy (anonymous):

I need a data set containing 9 numbers where the standard deviation is 1?

OpenStudy (amistre64):

1,2,3,4,5,6,7,8,9 maybe?

OpenStudy (amistre64):

you need a variance of 1 as well variance is the sum of the squares divided by n

OpenStudy (amistre64):

so the sum of the squares has to be 9 as well

OpenStudy (amistre64):

\[\sum_{1}^9x^2=9\]

OpenStudy (amistre64):

9 different numbers? or just a data set of 9 elements?

OpenStudy (anonymous):

There must be 9 numbers in the set where the standard deviation would be 1

OpenStudy (amistre64):

well, the sum of x^2 from 1 to 9 is 285; so id say divide that by 285

OpenStudy (amistre64):

\[\frac{1}{\sqrt{285}},\frac{2}{\sqrt{285}},\frac{3}{\sqrt{285}},\frac{4}{\sqrt{285}},\frac{5}{\sqrt{285}},\frac{6}{\sqrt{285}},\frac{7}{\sqrt{285}},\frac{8}{\sqrt{285}},\frac{9}{\sqrt{285}}\]

OpenStudy (amistre64):

still off

OpenStudy (anonymous):

X (X-M)2 (1 or -1)2 1 M= 1+1+1+1+1+1+1+1+1=9 (1/9) x 9 = 1 √1 = 1 X= number in data set M=mean of data set #s

OpenStudy (amistre64):

1 has no deviation

OpenStudy (loser66):

what is it? from which course?

OpenStudy (anonymous):

business math

OpenStudy (amistre64):

a set of 1s has no deviation; no variance

OpenStudy (amistre64):

0,1,1,1,1,-1,-1,-1,-1 is fine tho

OpenStudy (amistre64):

ugh, sample or population?

OpenStudy (anonymous):

the question just says List a data set in which the standard deviation is 1. it's for a discussion forum. Neither sample or population is stated

OpenStudy (amistre64):

well, sample deviation would be the 0,1,1,1,1,-1,-1,-1,-1 or you could pull out some z scores off the normal distribution curve

OpenStudy (anonymous):

Is there any way for population to have a standard deviation of 1 with 9 elements?

OpenStudy (amistre64):

yes, but its tricker

OpenStudy (amistre64):

assuming a mean of zero, you would need 9 values that sum their squares to 9

OpenStudy (amistre64):

\[\sum_{-4}^{4}x^2 = 60\]

OpenStudy (amistre64):

so each term is 1/sqrt(60)

OpenStudy (amistre64):

which is off since that would equate to 1 not 9

OpenStudy (anonymous):

Why wouldn't -1,-1,-1,-1,0,1,1,1,1 work?

OpenStudy (amistre64):

k sqrt(9/60), for k=-4 to 4 would work

OpenStudy (anonymous):

Has a sample standard deviation of 1

OpenStudy (amistre64):

that wont work for a population variance since 8/9 is not 1

OpenStudy (anonymous):

Are we using population or sample?

OpenStudy (amistre64):

either one is what they say, but they are now curious about pop var

OpenStudy (anonymous):

@vinnv226 neither was stated, but my guess is population.

OpenStudy (amistre64):

-4 sqrt(3/20) -3 sqrt(3/20) -2 sqrt(3/20) -1 sqrt(3/20) 0 sqrt(3/20) 1 sqrt(3/20) 2 sqrt(3/20) 3 sqrt(3/20) 4 sqrt(3/20)

OpenStudy (anonymous):

My statistics need some work, so I'll leave this to you guys

OpenStudy (anonymous):

ok, and to find a set of 9 elements that have a standard deviation of 0, wouldn't all 9 elements be the same number?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

if noting varies ... then there is no deviation

OpenStudy (anonymous):

Thank you

OpenStudy (amistre64):

youre welcome